OpenStudy (anonymous):

Show that the relation y^2 = 10x - x^2 can be expressed in polar coordinates as r = 10 cos theta. Help please!

OpenStudy (anonymous):

sigh...just think about it y = rsin(angle) x = rcos(angle) y^2+ x^2 = r^2 * (cos(angle)^2 + sin(angle)^2) = 10 * r cos(angle) => r^2 = 10 r cos angle => r = 10 cos angle

OpenStudy (anonymous):

Add the x^2 over. x^2+y^2=10x. Well x^2+y^2=r^2 and x=rcos(theta) So just plug in: r^2=10rcos(theta) r=10cos(theta)

OpenStudy (amistre64):

r^2 = x^2+y^2 x = r cos(t) x^2 +y^2 = 10x ; change to indentities r^2 = 10r cos(t) ; /r r = 10 cos(t)